Two bikes are 100 miles apart and traveling towards each other at 10 and 15 miles per hour. How long until they meet? A bee is traveling at 20 miles per hour and starts from the front tire of the slower bike and moves back and forth between the bikes without losing any speed. How far will the bee travel?
Answer by mananth(16946)

bikes are moving towards each other
so rate of approach = 10+15 = 25 mph
d= 100
t= d/r
100/25
t= 4 hours
the bee 's rate is 20 mph
it flies till they meet that is 4 hours
so it travels a distance of 20*4 = 80 miles